How do you change "x+7=0, y-1=0, y=1/4x+3, and y=-2/3x-1" into standard form
with integer coefficients?

Answer by edjones(7311)

General from: ax+by+c=0.
The 1st and 2nd can't be put in the general form because one of the coefficients is 0.
3)
y=x/4+3
y-y=x/4-y+3
x/4-y+3=0
4/4x-4y+3*4=0
x-4y+12=0
4)
y=-2/3x-1
2x/3+y+1=-2x/3-1+2x/3+1
2x/3+y+1=0
2x+3y+3=0
.
Ed